OpenStudy (anonymous):

what are the points to this y=8/7x+8

OpenStudy (anonymous):

..... Similary move, from (0,3) 8 units down the y axis and 7 units to the left.(-5,7) Mark these two points and draw the line.

OpenStudy (anonymous):

y=8/7x + 3 .. The y intercept point (0,3) .. slope = rise/run slope = 8/7 so from (0,3)move 8 units up the y axis and 7 units to the right